OpenStudy (anonymous):

if 7x+2y-6z = 12 ;x,y and z are positive what is the value of (2+z)/(7x+2y) A) 1/12 B) 1/6 C) 1/4 D) 5/12 E) 7/12

OpenStudy (phi):

notice what you get if you add +6z to both sides of the equation 7x + 2y -6z +6z=12+ 6z (7x+2y)= 12+6z can you factor 6 out of the expression on the right side of the equation?